Use the following to answer questions 5 and 6

On November 1, 20x1 Smarty Pants Company, a U.S. corporation, purchased diamonds from a Ukrainian company for 3,000,000 rubles, payable in 3 months. The relevant exchange rates between the U.S. and Russian currencies are given:

                                             Spot rate                      Forward rate (at February 1, 20x2)

November 1, 20x1               $0.456                                     $0.456

December 31, 20x1             $0.489                                     $0.481

February 1, 20x2                 $0.443

The company's incremental borrowing rate provides a discount rate of 0.945 for three months.

.

  1. If Smarty Pants does not attempt to hedge this transaction, what is the gain or loss that should be shown on the company's December 31, 20x1 financial statements?
  1. Assume that on November 1, 20x1 Smarty Pants Company enters a forward contract to buy 3,000,000 rubles on February 1, 20x2. How should Smarty Pants report the forward contract on December 31, 20x1?
Answer #1
1] Amount recorded as payable = 3000000/0.456 = $        65,78,947
Value of the payable on December 31, 20x1 = 3000000/0.489 = $        61,34,969
Gain on exchange rate movement $          4,43,978
2] Value of the forward contract on November 1 = 3000000/0.456 = $        65,78,947
Value of the forward contract on December 31 = 3000000/0.481 = $        62,37,006
Loss on forward contract $          3,41,941
The forward contract will be reported as a liability at $          3,41,941